# Rectified Iterative Disparity for Stereo Matching

### v1

@@ Abstract
I Introduction
@@
II Related Work
II-A Iterative-based Methods
@@
II-B Estimation of Disparity Uncertainty
@@ III Methods
III-A Cost volume-based Disparity Uncertainty Estimation
@@
where $a$ is the transition distance between correct to incorrect predictions. As the $a$ tends to infinity, Eq.6 is approximately equivalent to Eq.3. In Section IV-D, we explore the effect of the uncertainty ground truth setting on the accuracy of the uncertainty estimation and find that the best results are obtained with $a=1.5$ and $thr=3.0$.
III-B Disparity Rectification
@@
Notably, the UDC and UDR can be incorporated repeatedly into iteration-based methods at almost negligible cost (see Section IV-D for more details).
III-C Disparity Rectification Loss
@@
where $\alpha$ and $\beta$ are used to control the distribution and the overall scale of weights. It is shown in Section III-D how to combine this weight with L1 loss and Smooth L1 loss. In this paper, we set $\alpha$ to 1/8 and $\beta$ to 1/10, and demonstrate its validity experimentally (see Section IV-D and IV-F for more details).
III-D Loss Function
@@ IV-B Implementation Details
IV-C Estimation of Disparity Uncertainty
@@ IV-D Ablation Study
Combinations of proposed methods: As shown in Table V, we experiment with multiple combinations of the proposed methods. Comparing the results with those in Tables III and IX, our methods achieve further improvements after combination. In addition, we record the increase in the number of parameters, time consumption after inserting the design modules. It can be seen that inserting the UDR and UDC puts little burden for the model. As we present in Section III-B, the combination of UDR and UDC does not increase the number of parameters compared to a single method, since they share parameters from the UEC module.
IV-E Update on Bad Initial Disparity
@@ IV-F Benchmarks
IV-G Zero-shot Generalization
@@ VII Principle of UDC
VIII Other attempts of DR loss

### v2

@@ Abstract
I recommendation
Frankly speaking, this work serves as a preliminary exploration for SR-Stereo, and the paper writing requires some refinement. However, the incorporation of uncertainty in our approach may provide valuable insights for your research. I highly recommend reading the SR-Stereo paper [27] directly to gain a deeper understanding of the topic.
II Introduction
@@
III Related Work
III-A Iterative-based Methods
@@
III-B Estimation of Disparity Uncertainty
@@ IV Methods
IV-A Cost volume-based Disparity Uncertainty Estimation
@@
where $a$ is the transition distance between correct to incorrect predictions. As the $a$ tends to infinity, Eq.6 is approximately equivalent to Eq.3. In Section V-D, we explore the effect of the uncertainty ground truth setting on the accuracy of the uncertainty estimation and find that the best results are obtained with $a=1.5$ and $thr=3.0$.
IV-B Disparity Rectification
@@
Notably, the UDC and UDR can be incorporated repeatedly into iteration-based methods at almost negligible cost (see Section V-D for more details).
IV-C Disparity Rectification Loss
@@
where $\alpha$ and $\beta$ are used to control the distribution and the overall scale of weights. It is shown in Section IV-D how to combine this weight with L1 loss and Smooth L1 loss. In this paper, we set $\alpha$ to 1/8 and $\beta$ to 1/10, and demonstrate its validity experimentally (see Section V-D and V-F for more details).
IV-D Loss Function
@@ V-B Implementation Details
V-C Estimation of Disparity Uncertainty
@@ V-D Ablation Study
Combinations of proposed methods: As shown in Table V, we experiment with multiple combinations of the proposed methods. Comparing the results with those in Tables III and IX, our methods achieve further improvements after combination. In addition, we record the increase in the number of parameters, time consumption after inserting the design modules. It can be seen that inserting the UDR and UDC puts little burden for the model. As we present in Section IV-B, the combination of UDR and UDC does not increase the number of parameters compared to a single method, since they share parameters from the UEC module.
V-E Update on Bad Initial Disparity
@@ V-F Benchmarks
V-G Zero-shot Generalization
@@ VIII Principle of UDC
IX Other attempts of DR loss